OpenStudy (anonymous):

a = 1 + b, b = 5 − 2a

myininaya (myininaya):

so if a=1+b then b=5-2a=5-2(1+b)=5-2-2b=3-2b => b=3-2b you can solve b=3-2b for b?

myininaya (myininaya):

so we have b=3-2b add 2b on both sides 3b=3 => b=1

OpenStudy (anonymous):

i came up with a=1+b a+1=b

myininaya (myininaya):

if b=1 then a=1+b=1+1=2

myininaya (myininaya):

you check (a,b)=(2,1) this solution if you want by plugging it into the set of equations and see if each equations holds 2=1+1 1=5-2(2)

OpenStudy (anonymous):

what isn't right

OpenStudy (anonymous):

my bad. tried to go too fast. b=1 and a=2 myininaya is correct

OpenStudy (anonymous):

tks am just not good at this